# Point2RBox-v3: Self-Bootstrapping from Point Annotations via Integrated Pseudo-Label Refinement and Utilization

### v2

@@ 4 Experiments
3CS: CS denotes the abbreviation for Class-Specific Watershed. The Appendix A.6 contains detailed information on Class-Specific Watershed.
Table 4.1 compares Point2RBox-v3 with SOTA methods. These methods can be categorized into two tracks:
@@ 4.2 Results on More Datasets
RetinaNet (2017b) 68.69 60.57 47.00 54.96 21.80 57.67
GWD (2021) 71.66 63.27 48.87 57.60 25.30 57.80
FCOS (2019) 72.44 64.53 51.77 59.83 28.10 66.66
S2A-Net (2022) 75.81 66.53 52.39 61.41 27.30 66.47
@@
Sun et al. (2021) $\times$ 38.60 - - - - -
H2RBox (2023) 70.05 61.70 48.68 57.80 17.20 49.92
H2RBox-v2 (2023) 72.31 64.76 50.33 57.64 27.30 65.16
AFWS (2024) 72.55 65.92 51.73 59.07 - -
@@
P2RBox (2024) $\times$ 59.04 - - - - -
PMHO (2024b) $\times$ - - - 46.20 - -
PointOBB (2024) $\times$ 30.08 10.66 5.53 37.31 9.19 13.80
Point2RBox+SK (2024) 40.27 30.51 23.43 27.34 7.86 27.81
PointOBB-v2 (2025) $\times$ 41.68 30.59 20.64 39.56 9.00 18.99
PMS-SAM-RSD (2025) 46.00 - - - - -
PointOBB-v3 (2025b) 41.20 31.25 22.82 37.60 11.31 15.84
PointOBB-v3 (2025b) $\times$ 49.24 33.79 23.52 40.18 12.85 22.60
Point2RBox-v2 (2025a) 51.00 39.45 27.11 34.70 7.80 28.60
Point2RBox-v2 (2025a) $\times$ 62.61 54.06 38.79 44.45 14.20 30.90
[…]
* Comparison tracks: $\checkmark$ = End-to-end training and testing; $\times$ = Generating pseudo labels to train the FCOS detector (two-stage training). P2RBox/PMS-SAM-RSD/PMHO/Point2RBox-v3: Pre-trained SAM model; Point2RBox+SK: One-shot sketches for each class.
As shown in Table 4.2, Point2RBox-v3 also performs excellently on other datasets. On the more challenging DOTA-v1.5/2.0 datasets, Point2RBox-v3 shows a similar leading trend. In the end-to-end track, its AP50 is $17.09\%$ and $11.65\%$ higher than PointOBB-v3, respectively, and it consistently outperforms its predecessor, Point2RBox-v2 ($56.86\%$ / $41.28\%$ vs. $54.06\%$ / $38.79\%$). On the DIOR dataset, where object distribution is relatively sparse, our method achieves an AP50 of $46.40\%$. This result is slightly higher than PMHO ($46.20\%$), whose five-stage “point - proposal bag - mask - horizontal box - rotated box” pipeline makes it expensive to train and susceptible to error accumulation, while evidently higher than other methods. On the fine-grained dataset STAR, our method achieves a competitive AP50 of $19.60\%$. Furthermore, on the SAR image dataset RSAR, Point2RBox-v3 obtains an AP50 of $45.96\%$, demonstrating its effectiveness across different imaging modalities.
@@ 4.3 Ablation Studies
Tables 5-5 display the ablation studies on DOTA-v1.0. “E2E” denotes end-to-end training; “FCOS” denotes two-stage training (i.e. generating pseudo labels to train FCOS). The final values adopted are highlighted in gray.
@@
4.4 Application to Partial Weakly-Supervised Learning
Table 6: AP50 comparison on DOTA-v1.0/v1.5 under the partial weakly-supervised setting.“10%” means using only 10% point-labeled training data, the rest unlabeled.
@@ 4.4 Application to Partial Weakly-Supervised Learning
[…]
We integrated this method into the “Partially Weakly Supervised Oriented Object Detection”, i.e. PWOOD framework (Liu et al., 2025a), to demonstrate the universality and practicality of our approach. The operating principle of this framework is to train the model using a small portion of weakly labeled data (such as point annotations) and a large portion of unlabeled samples. We added our modules to the training process of the PWOOD framework and conducted experiments on the DOTA-v1.0 and DOTA-v1.5 datasets following the training process of PWOOD, setting the proportion of weakly labeled (point labeled) data to 10%, 20%, and 30% respectively.
@@ 5 Conclusion
ETHICS STATEMENT
This research adheres to the ICLR ethical guidelines and upholds the principles of responsible research. We ensure that no personally identifiable, sensitive, or harmful data were used. Our experiments were based on publicly available datasets and did not involve any human subjects or vulnerable groups. We have considered the potential societal impact of our methods, including the risk of misuse, and believe that these contributions primarily advance scientific understanding and do not pose foreseeable harm.
REPRODUCIBILITY STATEMENT
We follow the reproducibility guidelines in the ICLR 2026 author guidelines. We will open source code, configuration files, scripts and checkpoints to reproduce our results, including dataset construction, model training, and evaluation, on Github as soon as possible.
Appendix A Appendix
A.1 Loss Functions Inherited from Point2RBox-v2
In our Point2RBox-v3, we incorporate several loss functions that were originally proposed in Point2RBox-v2 (Yu et al., 2025a). These methods are not part of our novel contributions but are utilized to maintain a strong baseline performance. We group them under a single loss term, $\mathcal{L}_{\text{others}}$, which is a weighted sum of the following components.
A.1.1 Gaussian Overlap Loss ($\mathcal{L}_{O}$)
This loss function is designed to enforce mutual exclusivity among instances, which is particularly effective in densely packed scenes. As detailed in Point2RBox-v2, it operates by representing each oriented object as a 2D Gaussian distribution, $\mathcal{N}(\bm{\mu},\bm{\Sigma})$, and then minimizing the spatial overlap between these distributions. The overlap between any two distributions, $\mathcal{N}_{i}$ and $\mathcal{N}_{j}$, is quantified using the Bhattacharyya coefficient (Yang et al., 2021). For a given image with $N$ instances, an overlap matrix $\mathbf{M}\in\mathbb{R}^{N\times N}$ is constructed, where each element $M_{i,j}$ represents the coefficient between instance $i$ and $j$. The Gaussian overlap loss is then formulated as the sum of all off-diagonal elements, encouraging the detector to produce more compact and non-overlapping predictions:
$\mathcal{L}_{O}=\frac{1}{N}\sum_{i\neq j}M_{i,j}.$ (9)
A.1.2 Edge Loss ($\mathcal{L}_{E}$)
The Edge Loss is employed to refine the predicted bounding boxes by aligning their boundaries precisely with the object edges visible in the image. The process begins by applying an edge detection filter to the input image. For each predicted rotated bounding box (RBox), a corresponding feature region is extracted from the resulting edge map via Rotated RoI Align. By analyzing the edge intensity distribution within this region, new, more accurate regression targets for the box’s width ($w_{t}$) and height ($h_{t}$) are computed. The loss is then calculated using a smooth $L_{1}$ loss function, which penalizes the deviation between the predicted dimensions $(w,h)$ and the edge-aligned targets $(w_{t},h_{t})$:
$\mathcal{L}_{E}=\text{smooth}_{L1}([w,h],[w_{t},h_{t}]).$ (10)
A.1.3 Copy-paste Augmentation ($\mathcal{L}_{\text{box}}$)
To enhance the model’s robustness and generalization, especially in varied contexts, we utilize the copy-paste data augmentation strategy (Ghiasi et al., 2021), following its implementation in Point2RBox-v2. During the training process, object instances detected in a preceding step (e.g., step $k$) are cropped and subsequently pasted onto the training images for the current step ($k+1$). The bounding boxes of these pasted instances serve as additional ground-truth targets. The associated regression loss, which we denote as $\mathcal{L}_{\text{box}}$ and include in our $\mathcal{L}_{\text{others}}$ term, is calculated using the Gaussian Wasserstein Distance Loss (Yang et al., 2021).
A.1.4 Self-supervised Consistency Loss ($\mathcal{L}_{ss}$)
To further regularize the model in a self-supervised manner, we adopt the consistency loss introduced in Point2RBox-v2. This loss enforces prediction consistency between an original training image and its augmented counterpart (e.g., random rotation, flip, or scaling). Specifically, given an image $I$ and its transformed view $I_{\text{aug}}$, the detector outputs two sets of Gaussian representations and rotation angles:
$(\bm{\Sigma},\theta)=f_{\text{nn}}(I),\quad(\bm{\Sigma}_{\text{aug}},\theta_{\text{aug}})=f_{\text{nn}}(I_{\text{aug}}).$
The consistency loss is then computed as the sum of two terms: a Gaussian Wasserstein Distance loss that measures the discrepancy between the covariance matrices, and an angular regression loss that aligns the rotation predictions:
$\mathcal{L}_{ss}=\mathcal{L}_{\text{GWD}}(\alpha\bm{\Sigma}\alpha^{\top},\bm{\Sigma}_{\text{aug}})+\mathcal{L}_{\text{ANG}}(m\theta+R,\theta_{\text{aug}}),$ (11)
where $\alpha$ is the transformation matrix applied to $I$, $R$ is the rotation angle of augmentation, and $m\in\{1,-1\}$ accounts for symmetry in flips. This self-supervised term encourages the detector to capture consistent scale and orientation variations under different geometric transformations, thereby improving robustness.
A.2 Progressive Label Assignment (PLA)
Algorithm LABEL:alg:algorithm-pla delineates the detailed workflow of the PLA algorithm throughout the entire model training cycle.
$e$ is the number of current training epoch $E_{s}$ is the number of switch epoch $G$ is a set of ground-truth points on the image $V$ are the output Voronoi ridges (pixel coordinates) $I$ is the input image $S$ are the basin regions (pixel coordinates) corresponding to each annotated instance $PL$ is a set of pseudo labels on the image $L$ is the number of feature pyramid levels $A$ is a set of all anchor points $A_{i}$ is a set of anchor points on the $i$-th pyramid levels $a_{i}$ is a anchor point of $A_{i}$ Output: $P$ is a set of positive samples $N$ is a set of negative samples
1:if $e<E_{s}$ then
2:  V = Voronoi (G)
3:  S = Watershed (I, G, V)
4:  PL = minAreaRect (S)
5:  P, N = StandardLabelAssign (PL, A)
6:else
7:  $PL\leftarrow{}\varnothing$
8:  for each $g\in G$ do
9:   build an empty set for candidate pseudo label of $g$: $C_{g}\leftarrow{}\varnothing$
10:   for each level $i\in[1,L]$ do
11:    $S_{i}\leftarrow{}$ Select the prediction box with the closest $a_{i}$ to $g$ as the candidate
12:    $C_{g}=C_{g}\bigcup S_{i}$
13:   end for
14:   $PL_{g}=\arg\max_{b\in C_{g}}{score}(b)$
15:   $PL=PL\bigcup PL_{g}$
16:  end for
17:  P, N = StandardLabelAssign (PL, A)
18:end if
A.3 Analysis of Voronoi Watershed in Sparse Scenarios
The Voronoi Watershed method in Point2RBox-v2 (Yu et al., 2025a) constructs a Voronoi diagram. The annotation point of each object serves as a foreground marker for the watershed algorithm, and the background markers are defined by the Voronoi cell boundaries and regions where the fixed-size Gaussian distribution around the annotation point falls below a certain probability threshold. The Voronoi cells and Gaussian distributions are deeply related to the spatial layout of the objects.
This approach has a critical limitation in sparse scenes with minimal spatial layout information. The fixed size of the Gaussian distribution can lead to improperly sized foreground or background regions. This imbalance often causes the watershed algorithm to produce segmentation errors, including over-segmentation or under-segmentation. In contrast, the Segment Anything Model (SAM) (Kirillov et al., 2023) focuses primarily on object edges, textures, and other semantic visual cues. This makes it less dependent on the spatial layout of instances and allows it to produce more reliable segmentations in such challenging, sparse scenarios.
A.4 Detailed Formulation of Mask Selection Indicators
To select the optimal mask from the candidates generated by SAM, we use five indicators to score each mask based on simple prior knowledge provided by the user. This prior selection process is designed to be straightforward: the user only needs to make a simple judgment about the target’s general shape (e.g., more rectangular or circular) and the rough range of its aspect ratio, guided by example images from the dataset. The five indicators are center alignment, color consistency, rectangularity, circularity, and aspect ratio reliability. The setting of $w_{k,c_{j}}$ is as follows. When an object class has a significant corresponding geometric feature, the weight is set to a positive value. If the geometric feature is unimportant or ambiguous, it is set to 0. If the class does not possess the feature but is prone to being misinterpreted by a part that does, the weight is set to a negative value. For example, the weight for circularity for a basketball court is set to a negative value to penalize segmentations that only capture the center circle.
Briefly, center alignment, inspired by P2RBox (Cao et al., 2023), prioritizes masks whose centroids are close to the prompt point. Color consistency favors masks with a uniform color distribution, calculated based on the standard deviation of pixel values within the mask. Rectangularity and circularity measure how closely a mask’s shape resembles a standard rectangle or circle, respectively. Finally, aspect ratio reliability ensures the mask’s aspect ratio falls within a plausible, user-defined range. The mathematical formulation for each indicator is designed to be simple and computationally efficient.
1) Center alignment. This metric first checks if the prompt point is inside the mask’s minimum area bounding rectangle. If not, the mask is heavily penalized. If it is inside, the score $S_{align}$ is calculated based on the distance $d$ between the prompt point and the center of the rectangle using a Gaussian function.
$S_{align}=\exp\left(-\frac{d^{2}}{2\sigma_{c}^{2}}\right),$ (12)
where $\sigma_{c}$ is a scaling factor proportional to the diagonal of the image, ensuring that a smaller distance yields a higher score.
2) Color consistency. This metric evaluates color uniformity by calculating the weighted standard deviation of pixel values within the mask, giving higher weight to pixels near the prompt point. The final score $S_{color}$ is mapped from the average weighted standard deviation $\bar{\sigma}_{w}$ using an exponential decay function:
$S_{color}=\exp\left(-\frac{\bar{\sigma}_{w}}{\lambda}\right),$ (13)
where $\lambda$ is an empirical scaling constant. A lower deviation (more uniform color) results in a score closer to 1.
3) Rectangularity and Circularity. These metrics quantify how closely the mask’s shape resembles a rectangle or a circle. They are defined by the following ratios:
$S_{rect}=\frac{A_{mask}}{A_{mabr}},S_{circ}=\frac{A_{mask}}{A_{mcc}},$ (14)
where $A_{mask}$ is the mask area, while $A_{mabr}$ and $A_{mcc}$ are the areas of the minimum area bounding rectangle and the minimum circumscribed circle, respectively. To handle shapes near the edge of the image, the areas $A_{mabr}$ and $A_{mcc}$ are calculated from the parts of the shapes that lie within the image boundaries, that is, the part beyond the image boundaries is cropped and discarded.
4) Aspect ratio reliability. This metric checks if the aspect ratio of the mask’s minimum area bounding rectangle, $AR=\frac{\max(w,h)}{\min(w,h)}$, is within a predefined range $R_{min},R_{max}]$. If $AR$ is within this range, the score is 1. If it falls outside, the score $S_{ar}$ is calculated based on its deviation $D$ from the range:
$S_{ar}=\exp(-k\cdot D),$ (15)
where $k$ is a decay coefficient and the deviation $D$ is defined as $(R_{min}/AR-1)$ if $AR<R_{min}$ or $(AR/R_{max}-1)$ if $AR>R_{max}$. We found that for all non-extremely long-to-wide ratio categories in the remote sensing field, their long-to-wide ratio ranges were within the interval [1, 5], while the predicted incorrect masks were generally outside this range. For example, the fuselage of a plane (excluding the wings). Therefore, in order to enhance the usability of the model, we always set $R_{min},R_{max}]$ to be [1, 5].
A.5 Experimental Details
Figure 5: Efficacy of Class-Specific Watershed on Overlapping Instances. The top row shows challenging cases with overlapping objects. The standard Watershed algorithm produces unsatisfactory masks under these conditions (second row). Our class-specific Watershed (third and fourth rows) effectively mitigates this issue, leading to a noticeable improvement in mask separation and quality.
Implementation Details. Experiments are carried out using PyTorch 1.13.1 (Paszke et al., 2019) and the rotation detection toolkit MMRotate 1.0.0 (Zhou et al., 2022). All experiments follow the same hyper-parameters. We adopt Average Precision (AP) as the primary evaluation metric. All models are based on the ResNet50 (He et al., 2016) backbone and are trained using AdamW (Loshchilov & Hutter, 2018).
  1. 1.
  2. 2.
    Epochs: 12 for all datasets.
  3. 3.
    Augmentation: Random flip for all datasets.
  4. 4.
  5. 5.
    Multi-scale: All experiments are evaluated without multi-scale techniques (Zhou et al., 2022).
  • DOTA (Xia et al., 2018). DOTA-v1.0 contains 2,806 aerial images with 15 categories. DOTA-v1.5/2.0 are its extended versions with more small objects and new categories.
  • DIOR (Li et al., 2020). It is an aerial image dataset re-annotated with RBoxes based on its original HBox version (Cheng et al., 2022), featuring high variation in object scale and high intra-class diversity.
  • STAR (Li et al., 2024). An extensive dataset for scene graph generation, covering over 210,000 objects with diverse spatial resolutions and 48 fine-grained categories.
  • RSAR (Zhang et al., 2025c). A remote sensing dataset based on Synthetic Aperture Radar (SAR) imagery, containing 6 categories.
A.6 Watershed Algorithm vs Class-specific Watershed Algorithm
The quality of masks is critical for point-supervised learning. We observe that the mask quality produced by the Watershed algorithm severely degrades when processing overlapping objects, as illustrated in the second row of Figure 5. Motivated by this observation, we introduce a simple yet effective trick, termed class-specific Watershed. The key idea is to construct Voronoi diagrams and perform the Watershed algorithm on a per-class basis, rather than applying the algorithm indiscriminately to all objects in the image. Taking the ground track field (GTF) and soccer ball field (SBF) categories, which suffer from the most severe overlaps, as examples, the third and fourth rows of Figure 5 visually demonstrate the effectiveness of our class-specific Watershed. The results show a noticeable improvement in the mask quality for overlapping objects. This visual gain is corroborated by quantitative metrics; as indicated in Table 4.1 (see Point2RBox-v2+PLA+CS), this trick brings an overall performance gain of 1.45% (56.55% vs. 58.00%). Notably, the proposed trick not only ameliorates the segmentation of overlapping objects but also enhances the model’s overall performance. Class-specific Watershed and PGDM share a similar functional role. They can be considered as complementary strategies, and the optimal choice depends on the specific characteristics of the application scenario.
A.7 Limitation
The primary advantage of our Progressive Label Assignment (PLA) module—its adaptability to large scale variations—also defines its main limitation. The performance gain offered by PLA might be less pronounced on datasets where object scales are relatively uniform, as its benefits are most apparent in multi-scale scenarios. While introducing inductive biases like PLA into point-supervised task is effective, we believe that more fundamental advancements will likely stem from holistic architectural upgrades and more inventive loss function designs. Such directions could potentially elicit richer learning signals from extremely sparse annotations.
Furthermore, while our method alleviates the shortcomings of Point2RBox-v2 in sparse scenes by imposing spatial constraints, its performance is ultimately still limited by the characteristics of the existing SAM model. Since SAM is significantly more sensitive to color rather than texture and edges, the effectiveness of our method is limited in those sparse instances that blend with the surrounding environment or have unclear boundaries (such as basketball courts and ground track fields, see table 4.1), which led to its performance in some categories being even worse than that of previous models. In the future, researchers can consider enriching the prompts for SAM, such as introducing edge-aware algorithms to generate rough mask prompts, thereby improving performance in these challenging scenarios.
The Use of Large Language Models (LLMs)
In this study, the ideas, analysis, and conclusions presented are the sole product of the authors’ original thought and research. We utilized large language models (LLMs) to accomplish several auxiliary tasks in order to enhance the presentation quality and clarity of our manuscripts. Specifically, LLMs were employed for:
  • Table Generation: Create tables to reduce manual formatting workload.
  • Text Optimization: Simplify and refine sentence structures to make the expression clearer.
  • Consistency Check: Ensure uniformity in terms and style across all sections.
All the content generated by the large language model is carefully reviewed and verified by the author to ensure the technical accuracy and compliance with our scientific contributions.

### v3

@@ 4.1 Experimental Details
3CS: CS denotes the abbreviation for Class-Specific Watershed. The Appendix 6.6 contains detailed information on Class-Specific Watershed.
The best score is in bold and the second-best is in underline.
Table 1 compares Point2RBox-v3 with SOTA methods. These methods can be categorized into two tracks:
@@ 4.3 Results on More Datasets
RetinaNet [lin2017focal] 68.69 60.57 47.00 54.96 21.80 57.67
GWD [yang2021rethinking] 71.66 63.27 48.87 57.60 25.30 57.80
FCOS [tian2019fcos] 72.44 64.53 51.77 59.83 28.10 66.66
S2A-Net [han2022align] 75.81 66.53 52.39 61.41 27.30 66.47
@@
Sun et al. [sun2021oriented] $\times$ 38.60 - - - - -
H2RBox [yang2023h2rbox] 70.05 61.70 48.68 57.80 17.20 49.92
H2RBox-v2 [yu2023h2rboxv2] 72.31 64.76 50.33 57.64 27.30 65.16
AFWS [lu2024afws] 72.55 65.92 51.73 59.07 - -
@@
P2RBox [yu2024point2rbox] $\times$ 59.04 - - - - -
PMHO [zhang2024pmho] $\times$ - - - 46.20 - -
PointOBB [luo2024pointobb] $\times$ 30.08 10.66 5.53 37.31 9.19 13.80
Point2RBox+SK [yu2024point2rbox] 40.27 30.51 23.43 27.34 7.86 27.81
PointOBB-v2 [ren2024pointobbv2] $\times$ 41.68 30.59 20.64 39.56 9.00 18.99
PMS-SAM-RSD [lu2025semantic] 46.00 - - - - -
PointOBB-v3 [zhang2025pointobbv3] 41.20 31.25 22.82 37.60 11.31 15.84
PointOBB-v3 [zhang2025pointobbv3] $\times$ 49.24 33.79 23.52 40.18 12.85 22.60
Point2RBox-v2 [yu2025point2rbox] 51.00 39.45 27.11 34.70 7.80 28.60
Point2RBox-v2 [yu2025point2rbox] $\times$ 62.61 54.06 38.79 44.45 14.20 30.90
Point2RBox-v3 (ours) 59.61 48.34 34.47 41.50 14.60 40.80
Point2RBox-v3 (ours) $\times$ 66.09 56.86 41.28 46.40 19.60 45.96
* Comparison tracks: $\checkmark$ = End-to-end training and testing; $\times$ = Generating pseudo labels to train the FCOS detector (two-stage training). P2RBox/PMS-SAM-RSD/PMHO/Point2RBox-v3: Pre-trained SAM model; Point2RBox+SK: One-shot sketches for each class. The best score is in bold and the second-best is in underline.
As shown in Table 2, Point2RBox-v3 also performs excellently on other datasets. On the more challenging DOTA-v1.5/2.0 datasets, Point2RBox-v3 shows a similar leading trend. In the end-to-end track, its AP50 is $17.09\%$ and $11.65\%$ higher than PointOBB-v3, respectively, and it consistently outperforms its predecessor, Point2RBox-v2 ($56.86\%$ / $41.28\%$ vs. $54.06\%$ / $38.79\%$). On the DIOR dataset, where object distribution is relatively sparse, our method achieves an AP50 of $46.40\%$. This result is slightly higher than PMHO ($46.20\%$), whose five-stage “point - proposal bag - mask - horizontal box - rotated box” pipeline makes it expensive to train and susceptible to error accumulation, while evidently higher than other methods. On the fine-grained dataset STAR, our method achieves a competitive AP50 of $19.60\%$. Furthermore, on the SAR image dataset RSAR, Point2RBox-v3 obtains an AP50 of $45.96\%$, demonstrating its effectiveness across different imaging modalities.
@@ 4.4 Ablation Studies
Tables 5- 7 display the ablation studies on DOTA-v1.0. “E2E” denotes end-to-end training; “FCOS” denotes two-stage training (i.e. generating pseudo labels to train FCOS). The final values adopted are highlighted in gray.
@@
Necessity of Prior Knowledge in PGDM-Loss. It is crucial to note that the reliance on prior knowledge stems from the fact that SAM is trained on general-purpose datasets. We find that while SAM can effectively segment instances based on edge features in remote sensing scenes, its native confidence scores often fail to accurately reflect the quality of the generated masks in this specific domain. For example, in Figure 2, the ‘Prior-Guided Selector’ part shows that SAM will not tend to give the most correct answer a higher score. To validate this, we conducted an ablation study on DOTA-v1.0 as shown in Table 7. The results indicate that relying solely on SAM’s native confidence scores leads to a performance degradation, with AP50 dropping by 1.75 and 2.5 points for the end-to-end and two-stage frameworks, respectively. This empirical evidence strongly demonstrates the superiority of our proposed method.
Table 6: Ablation with Prior Knowledge in PGDM-Loss
Modules E2E FCOS
NoPrior-Loss 57.86 63.59
PGDM-Loss 59.61 66.09
Table 7: AP50 comparison on DOTA-v1.0/v1.5
@@
PLA + PGDM 50.67 52.22 58.02 64.57 43.25 47.05 50.51 56.18
We integrated this method into the “Partially Weakly Supervised Oriented Object Detection”, i.e. PWOOD framework [liu2025partial], to demonstrate the universality and practicality of our approach. The operating principle of this framework is to train the model using a small portion of weakly labeled data (such as point annotations) and a large portion of unlabeled samples. We added our modules to the training process of the PWOOD framework and conducted experiments on the DOTA-v1.0 and DOTA-v1.5 datasets following the training process of PWOOD, setting the proportion of weakly labeled (point labeled) data to 10%, 20%, and 30% respectively.